How many numbers are divisible by 2, 5, 9, and 13 between 100 and 100,000? How to implement this question in discrete math ?
1
Expert's answer
2020-11-27T17:23:15-0500

Since the numbers 2, 5, 9, and 13 are pairwise relatively prime, their the least common multiple is equal to "2\cdot 5\cdot 9\cdot 13=1170," and therefore, the number "n\in\mathbb N" is divisible by 2, 5, 9, and 13 if and only if "n" is divisible by "1,170."


The floor function "\lfloor x \rfloor" is defined to be the greatest integer less than or equal to the real number "x".


In discrete math is well-known the fact that the number of numbers that do not exceed "n\in \mathbb N" and are divisible by "d\in\mathbb N" is equal to "\lfloor \frac{n}{d} \rfloor". Since each number "n\le1000" is not divisible by 1170, between 1000 and 100,000 there are "\lfloor \frac{100,000}{1,170} \rfloor=85" numbers which are divisible by 2, 5, 9, and 13.


Answer: 85